### v1

# Localized Graph-Based Neural Dynamics Models for Terrain Manipulation

Abstract
Predictive models can be particularly helpful for robots to effectively manipulate terrains in construction sites and extraterrestrial surfaces. However, terrain state representations become extremely high-dimensional especially to capture fine-resolution details and when depth is unknown or unbounded. This paper introduces a learning-based approach for terrain dynamics modeling and manipulation, leveraging the Graph-based Neural Dynamics (GBND) framework to represent terrain deformation as motion of a graph of particles. Based on the principle that the moving portion of a terrain is usually localized, our approach builds a large terrain graph (potentially millions of particles) but only identifies a very small active subgraph (hundreds of particles) for predicting the outcomes of robot-terrain interaction. To minimize the size of the active subgraph we introduce a learning-based approach that identifies a small region of interest (RoI) based on the robot’s control inputs and the current scene. We also introduce a novel domain boundary feature encoding that allows GBNDs to perform accurate dynamics prediction in the RoI interior while avoiding particle penetration through RoI boundaries. Our proposed method is both orders of magnitude faster than naïve GBND and it achieves better overall prediction accuracy. We further evaluated our framework on excavation and shaping tasks on terrain with different granularity. The project page is available at chaoqi-liu.com/scoopbot.
I INTRODUCTION
Terrain manipulation is essential in construction industry and outer space exploration [1, 2]. Autonomous robots that can create desired structure in the construction site by interpreting sensory data and making decision on where and how to manipulate the terrain, can greatly improve efficiency and be deployed to place where lacks human labor, e.g., other planets. Prediction is a key challenge in terrain manipulation, since robots need to predict outcomes of excavated volume and the movement of terrain shape after scooping, dumping, or pushing material. People have traditionally built approximate models using analytical expressions and geometric calculations to predict excavation outcomes. However, these models are often limited in adapting to other action types and typically require predefined parameters [3]. Simulation-based methods use terramechanics physics simulators which can handle a wider variety of situations, but these are extremely computationally expensive and still require solving difficult parameter identification problems [4, 5, 6]. Learning-based approaches have also been investigated since early work in autonomous excavation [7], but recent attention has been focused on deep learning techniques, such as image-based prediction of granular media manipulation [8]. A line of recent work has studied the use of Graph-Based Neural Dynamics (GBND), a technique that uses Graph Neural Networks (GNNs) to model system dynamics of a scene represented as set of 3D particles, with each particle representing a small volume of matter [9, 10, 11]. GBND has been applied to various nontrivial deformable object manipulation tasks [12, 13, 14, 15], and can generalize well across different viewpoints, objects and robots poses, shapes and sizes. However, all prior such works have addressed table-top scenarios and objects with bounded volume that can be observed through static cameras.
Figure 1: Evaluation platforms. Our terrain manipulation system includes a UR5e robotic arm equipped with an overhead RGB-D camera. We are also extending this system to a mobile scooping platform for in-the-wild studies.
In this work, we propose a framework for modeling deformable terrains at fine resolutions over large spatial scales, such as a construction site in which a mobile excavator uses 3D mapping to generate a terrain representation. In particular, our method extends the GBND method to handle spatially large-scale and even unbounded terrains which may be represented by graphs of millions of particles that cannot fit in GPU memory. To predict the deformation of the terrain under a candidate interaction in a computationally efficient manner suitable for planning, our method restricts the attention of dynamics prediction to a dynamically-selected local neighborhood. We introduce a learning-based 3D region-of-interest (RoI) proposer that identifies a small subset of particles from the large graph that are likely to move. Assuming particles that are outside the RoI stay static during robot-terrain interaction improves speed by orders of magnitude in addition to higher dynamics prediction accuracy. Our method simultaneously learns the GBND dynamics and the RoI from simulation data, and both are fine-tuned on real data. Moreover, we introduce a novel boundary feature representation to address the problem that naïvely applying GBND to the identified subgraph leads to erroneous predictions. In our experiments, we verify both the efficiency and effectiveness of the proposed framework for manipulating terrains of different materials.
II BACKGROUND AND RELATED WORK
Terrain Modeling and Manipulation. Effective modeling of terrain movement induced by external forces, and its use in manipulation and locomotion, has been studied for years. Resistive Force Theory (RFT) and 3D-RFT, for example, is generic but reduced-order force models widely used in modeling locomotion in granular media [16]. However, RFT lacks the ability to model movements where inertial effects become significant, in addition to assumption of over-simplified geometry [7]. Other work instead use heightmap as the representation, and regress a latent linear dynamics model [17] for bulldozing. Linear dynamics models lack capability for more complicated interactions [8] and discretized heightmaps lack the flexibility of change of resolution, while the optimal resolution varies across tasks and underlying material [18]. In addition, particle-based representation provides high-fidelity geometry for tool equipped, focusing more on geometry during interaction, which provides extra precision. Alternatively, model-free methods are utilized widely [19], but these approaches often grapple with high sample complexity and exhibit a task-specific nature.
Dynamics Model Learning. Physics-based models have shown effectiveness in a wide range of robotic manipulation tasks [20, 21, 22], but accurate model building requires full state estimation. To address this issue, recent advances in machine learning are applied to learn the system dynamics models directly from sensory data [23, 24, 25, 26]. The learned dynamics model generally demonstrates superior effectiveness relative to its model-free counterparts [27]. Graph-based scene representation processed by GNNs have showcased effectiveness in dynamics modeling due to the spatial relational inductive bias injected [28, 29, 9, 10, 11]. Prior work explored a wide range of applications of graph-based dynamics model, including but not limited to rigid bodies [9, 10], elastic-plastic materials [13, 12], fluids [9, 11], and granular materials [18, 11, 30]. However, the vast majority of prior work model systems that are either confined to a 3D box or atop a 2D plane, and our work demonstrates an approach to extend graph-based approaches to terrains of unbounded breadth and depth using regions of interest and RoI boundary features.
Figure 2: System overview. The proposed dynamics model is shown in the central block: scene particles and the control input are fed into the RoI proposer, which selects particles likely to move based on the interaction caused by the control input. This is followed by a rollout using GBND. When integrated with planning, multiple trajectories are sampled and rolled out in parallel, and the best trajectory, determined by a pre-defined metric, is executed.
III METHOD
III-A Overview
Our goal is to learn a computationally-efficient model to predict the shape of a terrain after robot-terrain interaction, such that the model can be used in terrain manipulation planning. Specifically, letting $x_{t}$ denote the state of the terrain, we model $\hat{x}_{t+1}=f(x_{t},u_{t})$ where $u_{t}$ is the robot control and $t$ is the time index. We learn fine-grained prediction where the time step is 0.1 sec and $u_{t}$ indicates a small robot end effector displacement. The dynamics model is trained on data recorded during three types of actions (scooping, pushing, and dumping) and we use a unified dynamics model for all actions. Pretraining is performed from simulation data and then the model is fine-tuned on real data.
We choose GBND as our terrain model, which is a particle-based scene representation that provides flexibility in inter-particle interactions, an inductive bias toward locality in interactions, mass conservation, and preserves fine details of deforming solids and fluids [9, 11, 10].
The core challenge addressed by our method is that the terrain is a volume of material with a very high-dimensional state. Our approach assumes that only a small region of interest (RoI) $R_{t}(u_{t})\subset\mathbb{R}^{3}$ of the volume is affected by the interaction. In other words, letting $p_{t}^{(i)}$ denote the $i^{\text{th}}$ particle in the terrain at time $t$, its future state $\hat{p}_{t+1}^{(i)}=p_{t}^{(i)}$ for any point $p_{t}^{(i)}\notin R_{t}(u_{t})$. Our method simultaneously learns $f$ and $R_{t}$. The RoI proposer is formulated as a super-level set of a learned function $g(p,u_{t})$ such that $g(p,u_{t})\geq 0$ if and only if $p\in R_{t}(u_{t})$. Using the RoI, we only predict the motion of a subgraph of the particle graph. Then, given a terrain manipulation goal, a planner is used to find a near-optimal action sequence. The proposed framework is shown in Fig. 2.
Our key innovations within this framework include: 1) learning $f$ and $g$ to minimize GBND memory usage and computation time while retaining high accuracy; and 2) introducing novel node features that enable learning GBND over arbitrary volumetric domains, ensuring that accuracy is preserved at the boundary of $R_{t}$, and that translation invariance is maintained.
III-B Background on Graph Based Neural Dynamics
This section presents background on the GBND formulation, and for more details we refer readers to Refs. [10, 11, 31]. A GBND describes the environment (robot and terrain) as a set of interacting particles, each of which implicitly represents a small volume of matter, and interactions between particles are learned. The dynamics model is a GNN that represents the environment state at time $t$ as a directed graph (digraph) $x_{t}\equiv G_{t}(V_{t},E_{t})$, where $V_{t}\triangleq\{v_{t}^{(i)}\}_{i=1}^{|V_{t}|}$ is the set of nodes, and $E_{t}\triangleq\{e_{t}^{(k)}\}_{k=1}^{|E_{t}|}$ is the set of edges, with each edge $e_{t}^{(k)}$ denoting a potential direct interaction from $v_{t}^{(i)}$ to $v_{t}^{(j)}$. Indirect interactions between particles are modeled in this framework using a message passing algorithm.
In graph construction, each particle $p_{t}^{(i)}$ is instantiated into a node with feature vector $v_{t}^{(i)}$, and nearby particles are instantiated into edges with the feature vector $e_{t}^{(k)}$. The dynamics are defined as follows. Each node and edge are first encoded into a latent representation,
$z_{0}[v_{t}^{(i)}]=f_{V}^{\text{enc}}(v_{t}^{(i)}),\quad z_{0}[e_{t}^{(k)}]=f_{E}^{\text{enc}}(e_{t}^{(k)}).$ (1)
Then, message propagation is run for $L$ iterations to propagate, aggregate, and update latent representations across the graph. Mathematically, for $l\in\{0,1,...,L-1\}$,
$z_{l+1}[e_{t}^{(k)}]=f_{E}^{\text{prop}}(z_{l}[e_{t}^{(k)}],z_{l}^{i},z_{l}^{j})$ (2)
for an edge $e_{t}^{(k)}\triangleq(i,j)$, and messages along edges to nodes are propagated and aggregated via
$z_{l+1}[v_{t}^{(i)}]=f_{V}^{\text{prop}}\Big{(}z_{l}[v_{t}^{(i)}],\sum_{e\in E^{+}(v_{t}^{(i)})}z_{l}[e]\Big{)}.$ (3)
Finally, a node decoder $f_{V}^{\text{dec}}$ predicts displacement of the particle, $\Delta\hat{p}_{t}^{(i)}=f_{V}^{\text{dec}}(z_{L}[v_{t}^{(i)}])$. The ultimate dynamics prediction is given by $\hat{p}_{t+1}^{(i)}=p_{t}^{(i)}+\Delta\hat{p}_{t}^{(i)}$. GBND learns encoder, decoder, and propagation networks $f_{V}^{\text{enc}}$, $f_{E}^{\text{enc}}$, $f_{V}^{\text{dec}}$, $f_{V}^{\text{prop}}$, $f_{E}^{\text{prop}}$ via backpropagation through message passing.
III-C GBND for Large-Scale Terrain Modeling
We now describe how we use GBND to model terrains with high accuracy and low computational cost. A large-scale graph is first constructed through perception input (Fig. 3). For an observed terrain surface map, we instantiate terrain particles $P_{t}=\{p_{t}^{(i)}\}_{i=1}^{N_{p}}$, through the volume of the terrain down to a sufficiently large depth. We also instantiate tool particles sampled on the robot’s end effector $T_{t}=\{p_{t}^{(i)}\}_{i=N_{p}+1}^{N_{p}+N_{t}}$. For each particle, we connect it to the $k$ closest particles that meet a distance threshold through a nearest-neighbor algorithm.
Figure 3: Visualization of a typical scooping simulation. Our method (left) only predicts the dynamics of the highlighted particles in the predicted RoI. Predictions are computed for an order of magnitude fewer particles as the full graph (right) while retaining similar accuracy.
Our baseline node feature implementation is defined as
$v_{t}^{(i)}\triangleq\left[\mathsf{v}_{t-H:t-1}^{(i)},u_{t}^{(i)},c^{(i)}\right]^{\top},$ (4)
where $\mathsf{v}_{t-H:t-1}$ is particle $p^{(i)}$’s velocity over the last $H$ time steps, $u_{t}$ is the impulse on the particle directly caused by robot action, and $c$ is the particle class (terrain or tool) embedding. The edge features for an edge $e_{t}^{(k)}$ connecting particles $i$ and $j$ are
$e_{t}^{(k)}\triangleq\left[p_{t-H:t}^{(i)}-p_{t-H:t}^{(j)},c^{(i)},c^{(j)}\right]^{\top}.$ (5)
We will show in the next section how including boundary features with nodes $v_{t}^{(i)}$ leads to higher accuracy and generalization when predicting over an RoI.
Because the computation time and memory usage of GBND are closely tied to the number of particles and edges in the graph, limiting the graph’s size helps ensure efficiency. So, at each time step $t$, our RoI proposal method defines the volume $R_{t}(u_{t})\subset\mathbb{R}^{3}$ via the implicit function $g(p,u_{t})$. We describe details of our implementation in Section III-E. The terrain particles $p_{t}^{(i)}\in R_{t}(u_{t})$ are considered to be dynamic, so we extract the subgraph corresponding to these particles and then connect them to tool particles. Only the terrain and tool particles in this subgraph undergo node/edge feature instantiation and GBND propagation.
III-D Encoding RoI Boundaries as Node Features
The standard GBND machinery is best suited for tasks with similar training / test distributions, but our setting introduces the complication of unboundedness and varying regions of interest. As an example, past works using GBND for fluid and pile manipulation have used a fixed domain of particles, such that the network could implicitly learn the boundary conditions of a floor and walls [13, 12, 32, 18]. In our case, “boundary conditions” are not strictly-enforced boundaries, but are rather just more material that is unaffected by the current interactions due to energy dissipation.
In effect, particles adjacent to the RoI boundary should “feel” resistance as though the boundary were fixed. With baseline node encoding, these particles will simply pass through the boundary as if there was nothing there. We also experimented with adding “wall particles”, a boundary set of nodes, directly connected to nodes in the RoI, whose latent features are frozen. Not only does this increase memory usage due to massive wall particles, but in our experiments, inner particles were able to penetrate gaps in the sampled wall particles which introduces unnatural artifacts.
Our approach incorporates an estimated geometric normal of each particle within RoI. Normal estimation is performed using standard techniques within the point cloud of RoI particles [33]. In the interior, the normal information is not likely to help much, but normals are likely to be strongly aligned near the boundary. This allows our network to encode the effect that nodes near the RoI boundary should “feel” resistance in the direction of the normal.
This encoding also has the favorable effect of translational invariance, which is that for the overall dynamics model $f$,
$\forall\Delta p\in\mathbb{R}^{3},f(p_{t}^{(i)}+\Delta p,u_{t})=f(p_{t}^{(i)},u_{t})+\Delta p.$ (6)
This ensures that the model’s predictions are independent of the absolute positions of particles.
Figure 4: Comparing different methods for encoding RoI boundaries in node features. The baseline (no encoding) does not perform well, whereas encoding absolute height ($z$ features) leads to implicitly learning a “floor effect.” This becomes detrimental when the validation set has different heights from the training set. In contrast, our proposed normal features perform well in all cases.
Experiments in Fig. 4 present validation losses of learned GBND models for different sets of node features. As baselines we compare the standard GBND encoding as well as augmenting with particle height ($z$ features).In the in distribution case, the training and validation sets used terrains whose particles were sampled in a fixed $z$ range of [0 cm, 5 cm]. In the out of distribution case, the validation particle sets were randomly shifted in the $z$ axis by randomly sampled offsets from -100 to 100 m. Encoding the $z$ value significantly improves loss in the in-distribution case, demonstrating that the network is implicitly learning a “floor effect”. This implicit relationship is detrimental in the out-of-distribution case. In contrast, our proposed normal encoding leads to better generalization than the baseline both in and out of distribution.
III-E Convolutional Neural Net RoI Proposer
Finally, we describe how we encode and learn our RoI implicit function $g(p,u_{t})$. We first extract the planar tool position and yaw from $u_{t}$ and define a 2D reference frame $T_{\text{tool}}$. We then predict a tool-centered 2D height function $\tilde{g}(x,y)$ such that $g(p,u_{t})=p_{z}-\tilde{g}(T_{\text{tool}}^{-1}(p_{x},p_{y}))$. In other words, points whose $z$ exceeds $\tilde{g}(x,y)$ (in the tool-centered frame) are included in the RoI.
Our approach uses a terrain heightmap and instantiates $\tilde{g}$ as a Convolutional Neural Net (CNN) architecture. A heightmap is used because it is compatible with elevation mapping pipelines and avoids computations over the large number of volumetric particles. The heightmap with grid resolution 2.5 cm is centered about the tool and cropped to a wide enough range (set to 0.4 m $\times$ 0.4 m in our current implementation).
At each time step, particles in the bounding box are squeezed into a 32$\times$32 heightmap, with each pixel is concatenated with a 16D latent representation of the current control input. Our CNN is a U-net-like [34] architecture consisting of three levels of pooling, nonlinear, and upsampling steps with skip layers. The output CNN is a 16$\times$16 depth map representation of $\tilde{g}$.
Algorithm 1 Dynamics with Scale Reduction
procedure Propose($\{p_{t-H:t}^{(i)}\}_{i=1}^{N},u_{t}$)
    $M_{t-H:t}\leftarrow$ Project2D( $\{p_{t-H:t}^{(i)}\}_{i=1}^{N}$ )
    Predict $\tilde{g}$ based on $M_{t-H:t},u_{t}$triangleright $R_{t}$ in Sec. III-E
    for $i\in[N]$ do
         if $p_{t}^{(i)}.z\geq\tilde{g}(p_{t}^{(i)}.x\,,\,p_{t}^{(i)}.y)$ then
             $ndices ← i$
    return indices
procedure Step( $x_{t-H:t},u_{t}$)
[…]
    Align coordinate frame $L$ with scoop-tip
[…]
    Rollout dynamics $\forall i,$
[…]
    return $\{{}^{W}\hat{p}_{t+1}^{(i)}\}_{i=1}^{N}$
We train the RoI predictor with a simulation dataset. To identify particles that are moving, we first apply a velocity threshold of 10 cm/s. Noise in the simulator remains a challenge, and we employ connected component filtering to focus on the main bulk of the terrain movement near the scoop tip. The CNN is trained given the particle heightmap as input and the output is a heightmap representation of $g$. We define the target heightmap as a function that separates the moving and non-moving particles. This is defined by first taking the upper heightmap of non-moving particles, then creating a bowl-shaped function by inflating their value proportionally to the distance to the center, which we found empirically encourages training convergence. Then, we take the min of this map with the bottom-up heightmap of the moving particles. For grid cells in the particle heightmap that are non-empty, we apply an $L_{1}$ loss between the CNN output and the target heightmap. For empty grid cells, we set a hinge loss at a target value of 0.5 m, and do not penalize exceeding this value. The pseudocode of GBND coupled with scale reduction is presented in Algo. 1.
III-F Planning Terrain Manipulation Trajectories
Our planner chooses amongst parameterized scooping, pushing, and dumping trajectories to greedily minimize the difference of the predicted terrain shape to a target heightmap. After every trajectory is executed the large-scale terrain shape is updated and the planner is called again. For scooping and pushing, we inherit the penetrate-drag-scoop (PDS) trajectory parametrization widely used in excavation community [1, 35, 7] (Fig. 5), where the scoop maintains 0 roll angle throughout the entire trajectory. This defines a 9D motion parameter space. For dumping, we use a 6D parameter space for the dumping location, yaw angle, pre-dump and post-dump pitch angles. We also maintain the number of in-scoop particles, which is 0 after dump and push. Scooping and pushing are available if the scoop contains no material.
Figure 5: Sideview of a “scoop” primitive. The “push” primitive is a special case of “scoop” primitive with $\alpha=\beta=\pi/2$. The “dump” primitive is simply pivoting the scoop about the scoop-tip. Figure borrowed from [1].
Algorithm 2 Planning with Learned Dynamics
score $S$, scene $x_{t}$, goal description $\mathsf{G}$, motion parameter space $\Omega$, batch size $b$, time budget $\mathsf{B}$
$u^{*}_{0:T}\leftarrow nil$
while time used < $\mathsf{B}$  do
    Sample a batch of trajectory parameters $\omega_{1},\ldots,\omega_{b}$
    Convert batch of $\omega_{i}$ to control sequences $u_{i,0:T_{i}}$, $\forall i$
    Predict batch $\hat{x}_{i,t+T_{i}}\leftarrow$ Rollout($x_{t},u_{i,0:T_{i}}$), $\forall i$
    Evaluate scores $S(\hat{x}_{i,t+T_{i}},\mathsf{G})$, $\forall i$
    Update best control sequence $u^{*}_{0:T}$
return $u^{*}_{0:T}$
The optimization is based on the sampling-based planner MPPI [36], which uses the scores of prior trajectories to update a sampling distribution over the parameter space $\Omega$. In each iteration, a batch of motion parameters is sampled $\{\omega_{i}\}_{i=1}^{b}$, and these are then converted into corresponding control sequences, $\{u_{i,0:T_{i}}\}_{i=1}^{b}$. For all sequences, the dynamics model is rolled out (in parallel), and scores are evaluated. The planner maintains the best sequence and updates the distribution to focus on areas of the space in which the score is low. Pseudocode is listed in (Alg. 2).
IV EXPERIMENTS AND RESULTS
The computational performance and accuracy of the proposed method were thoroughly evaluated and contrasted against simpler baselines. Furthermore, the capability of our method was assessed through its application to extended-range excavation tasks. All experimental evaluations were conducted using an Intel i9-13900K CPU and a single NVIDIA GeForce RTX 4080 GPU.
IV-A Implementation Details
The architectures of GBND are MLPs, trained over 200,000 backpropagation steps, unless stated otherwise. For data collection during pre-training, the SAPIEN simulator [37] was utilized. Attempts to employ Nvidia Flex [38] were made, but it was abandoned due to the increasing prevalence of unrealistic behaviors as the particle count escalated. In the terrain generation phase, the diamond-square algorithm [39] was employed to produce landscapes comprised of approximately 3,000 miniature cubes. Each cube was then populated with 5 to 10 particles, and further reduction of terrain particle sets was achieved through farthest-point sampling [40].
Approximately 1,000 robot-terrain interaction trajectories, each uniquely parameterized, were simulated. Throughout these simulations, the downsampled terrain particles were meticulously tracked to facilitate the learning of the dynamics model. Training on this simulation data involved regressing the motion trajectories of the particles using the Mean-Square Error (MSE) criterion, aided by the perfect association of scene particles. Concurrently, the CNN proposer was trained as detailed in Section III-E. The model’s forward simulation steps were set at 0.1 seconds.
IV-B Computational Efficiency and Accuracy
Figure 6: Comparing our method against full-scale GBND (FS) and geometric region proposers with different size (Geo-X). Our method demonstrates significant advantages in speed and GPU memory. $\approx$ 3,000 particles and batch size 256 are used in these experiments, per sample measurements are reported (i.e., divided by 256). Colors and labels are shared by both figures.
The computational performance and accuracy of our method are compared against two baselines. The first baseline utilizes all particles within a large-scale terrain, while the second employs a naïve geometric Region of Interest (RoI), specifically, a cylinder centered at the scoop tip with a predetermined radius. To maintain computational feasibility with the full-terrain baseline, terrains are restricted to 3,000 particles — significantly fewer than what might be encountered by an excavator operating on an large worksite. Figure 6 presents the results on a terrain measuring 50 cm $\times$ 50 cm $\times$ 5 cm, with a particle density of 0.5 particles per cm3. Here, ‘FS’ denotes the full-scale terrain predictor, and ‘Geo-X’ represents the geometric RoI with a cylinder diameter of X cm. Each RoI-based approach is developed using the methodology described in Section III-E alongside our novel node boundary features.
In terms of per-sample processing time and GPU memory utilization, our method matches the speed of the smallest geometric RoI and requires even less GPU memory, while delivering superior accuracy. It is important to note that enhanced RoI prediction significantly improves the accuracy of dynamics predictions. This is because the GBND is prone to inaccuracies, such as predicting minimal movements of static particles. An RoI introduces a strong inductive bias that is absent in full-scale GBND, leading to jitter among distant particles. For smaller values of X, Geo-X exhibits large errors due to the exclusion of moving particles. Conversely, as X increases, the inclusion of numerous static particles leads to performance degradation due to GBND jitter.
Batch Ours Geo-6 Geo-12 Geo-18 FS SAPIEN
1 4.624 3.543 4.459 5.157 5.443 3.309 (CPU)
2 2.454 1.913 2.280 2.616 2.775 2.104
4 1.400 1.036 1.250 1.463 1.642 1.627
8 0.860 0.586 0.755 0.931 1.097 1.563
16 0.522 0.366 0.501 0.650 0.872 1.713
32 0.337 0.242 0.364 0.520 0.885 1.991
64 0.232 0.181 0.332 0.498 0.809 2.211
128 0.179 0.169 0.318 0.456 0.774 1.755
256 0.162 0.163 0.301 0.438 0.750 OoM
512 0.181 0.181 0.293 0.431 OoM OoM
1024 0.198 0.198 0.294 OoM OoM OoM
TABLE I: Average prediction time (ms) per sample as the batch size varies, showing that 256 is an optimal batch size for our framework. For SAPIEN, 3000 cubes are simulated.
Furthermore, since the planning speed and quality are heavily influenced by the batch size, memory efficiency becomes a critical objective for our dynamics model. Table I illustrates that our method reaches optimal performance at a batch size of $b=256$, achieving an average processing time of less than 0.2 ms per sample. Under these conditions, the planner is capable of rolling out hundreds of trajectories across numerous time steps in seconds. We also compare against the recently updated GPU-based simulator in SAPIEN 3.0 [37] in scenes with 3,000 cube-shaped particles, without a simulated robot. Our method is an order of magnitude faster and can handle larger batch sizes within available GPU memory. More importantly, it also allows fine-tuning to match real-world material behavior.
IV-C Real World Terrain Manipulation
Real-world experiments were conducted using the platform shown in Fig. 1, left, on two different materials as depicted in Fig. 7: (1) pebbles with an approximate grain size of 0.8 - 1.0 cm, and (2) play sand with a grain size significantly less than 1 mm. To ensure experimental repeatability, the terrain was contained within a box measuring 0.9 m $\times$ 0.6 m $\times$ 0.2 m, though the methodologies developed are equally applicable to natural terrains.
((a)) Pebbles 0.8 - 1.0 cm
((b)) Sand $\ll$ 1 mm
Figure 7: Materials used in real world experiments, U.S. quarter coin on the bottom right for scale. Figure borrowed from [1].
To adapt our model to real-world data, we sampled 100 parametrized robot-terrain interaction trajectories, monitored via an overhead RGB-D camera. Scene point clouds were derived from depth images, with robot proprioception utilized to exclude robot particles while selectively retaining in-tool (scoop) particles by re-sampling within the tool’s convex hull. Due to the inability to track individual particle trajectories across frames, measures such as the weighted sum of chamfer distance [41] and earth-mover’s distance [42] were employed. This fine-tuning process typically improves the dynamics models’ prediction accuracy by approximately 5–10%, compared to the simulation-trained model (Fig. 8).
((a)) Rock
((b)) Sand
Figure 8: Fine-tune (FT) the pre-trained (PT) GBND model on real-world datasets significantly improves the model stability and precision in practical scenarios, as evidenced by a reduced median error and a more compact interquartile range.
The experimental setup focused on two specific terrain shaping tasks: (1) creating a hole, and (2) forming a rectangular moat. The desired terrain configurations were represented as 2D heightmaps, with the $L_{1}$ loss between the current scene heightmap and the target heightmap serving as the performance metric. These target shapes have sharp discontinuities and cannot be physically realized due to the material’s settling behavior (angle of repose is approximately 25-40). For each manipulation task, our planner evaluated 1500 trajectories with an average horizon of 30 steps (equivalent to 3 seconds), requiring 20-30 seconds of planning time. Qualitative results are shown in Fig. 9, demonstrating successful terrain shaping.
Figure 9: Sequences of heightmaps recorded during real-world manipulation.
((a)) Hole Pebble
((b)) Hole Sand
((c)) Moat Pebble
((d)) Moat Sand
Figure 10: Errors between the observed terrain and the target terrain after each PDS trajectory. Interquartile ranges are shaded. 10 trials per scenario.
For a qualitative comparison, we implement a heuristic baseline that employs dynamic programming to determine the scooping trajectory that maximizes the intersection of the swept volume and the excess material in the current terrain heightmap, similar to the volume maximization approach developed by Yang et al. [43]. We also constrain the solver to use axis-aligned trajectories to boost its computational efficiency. Fig. 10 illustrates that the heuristic makes rapid initial progress but is unable to shape the terrain’s finer details by anticipating settling behavior.
V CONCLUSION AND LIMITATION
We presented a framework for handling unbounded terrains in graph-based neural dynamics (GBND) models that simultaneously learns a RoI and particle dynamics. Novel node features are key to success of this model. Experiments show that this method achieves orders of magnitude faster prediction and GPU memory usage compared with naïve GBND models or GPU-based physics simulators. It can be trained with both simulated and real data, and when used as a model for planning our method can achieve desired target shapes on real-world excavation platforms.
There remains some limitations of our work that we would like to address in future work. First, our RoI proposer does not accurately model long-range terrain effects (e.g., landslides). Second, our current platform uses an overhead camera to update the terrain, and we would like to extend this to egocentric 3D mapping, in which our model may need to model occluded parts of the terrain. Third, our models assume terrain homogeneity, and future work should address heterogeneous materials including terrains with embedded obstacles. Lastly, the current dynamics model requires full-parameter fine-tuning when moving to unseen materials, one potential solution is to adapt physics-parameter conditioned GBND [44] for online adaptation.